OpenStudy (ttp):

Evaluate 80 × a to the power of -2 for a = 4. 20 5 one-twentieth one-fifth

OpenStudy (anonymous):

5

OpenStudy (anonymous):

plug in 4 for a ..... so 80 times 4 ^ -2 using order of operations ....you would do exponents first .... so 4 ^-2 is what

OpenStudy (the_fizicx99):

80 * 4^-2 80 * 1/4^2 = 5

OpenStudy (anonymous):

\[80 \times a ^{-2} =80/a ^{2}=80/4^{2}\]

OpenStudy (anonymous):

please do not give answers ty and have a good day

OpenStudy (the_fizicx99):